a tractor with a purchase priece of 75000 INR and has to be used for plowing with a 3 bottom MB plough with a purchase priece of 60000 INR with suitable asssumption calculate the total cost of operation per hour with straight line method of depriciation


1
Expert's answer
2021-01-26T02:59:00-0500

Depreciation (D) = (P - S) / LH

P - price of tractor; S - price of plough;

L - life of tractor (15 years); H - annual use of tractor (1000 h)

D = (75000 - 60000) / (15 x 1000) = 1.0/h

Interest (I) = ((P + S) / 2) x (I / H)

I - interest rate (15%)

I = ((75000 + 60000) / 2) x (0.15 / 1000) = 10.125/h

Tax, Shelter & Insurance = (2% of purchase price)/1000 = 0.02 x 75000/1000 = 1.5/h

Total fixed cost = 1.0 + 10.125 + 1.5 = 12.625/h